### v1

# QIANets: Quantum-Integrated Adaptive Networks for Reduced Latency and Improved Inference Times in CNN Models

Abstract
Convolutional neural networks (CNNs) have made significant advances in computer vision tasks, yet their high inference times and latency often limit real-world applicability. While model compression techniques have gained popularity as solutions, they often overlook the critical balance between low latency and uncompromised accuracy. By harnessing quantum-inspired pruning, tensor decomposition and annealing-based matrix factorization – three quantum-inspired concepts – we introduce QIANets: a novel approach of redesigning the traditional GoogLeNet, DenseNet, and ResNet-18 model architectures to process more parameters and computations whilst maintaining low inference times. Despite experimental limitations, the method was tested and evaluated, demonstrating reductions in inference times, along with effective accuracy preservations.
1 Introduction
The field of computer vision (CV) has recently experienced a substantial rise in interest (Su & Crandall, 2021). This surge has created transformative advancements, driving the development of deep learning models, particularly those based on the convolutional architecture, such as DenseNet, GoogLeNet, and ResNet-18. These methods have significantly optimized neural networks for image processing tasks, achieving state-of-the-art performance across multiple benchmarks. (Anumol, 2023; He et al., 2015; Szegedy et al. 2015) However, the increasing computational complexity, memory consumption, and model size – often comprising millions to billions of parameters – pose substantial challenges for deployment, especially in time-sensitive and computationally-limited scenarios. The demand for low-latency processing in real-time applications, such as image processing and automated CV systems, is critical; compact models are needed for faster responses (Honegger et al., 2014).
To address these issues, researchers have explored various optimization techniques to reduce inference times and latency while maintaining high accuracy. Model compression techniques such as pruning, quantization, and knowledge distillation have shown promise in enhancing model efficiency (Li et al., 2023). Yet, these methods often come with trade-offs that can impact model performance, necessitating a careful balance between energy efficiency and accuracy.
In recent years, the principles of quantum computing have emerged as an avenue for accelerating inference in machine learning (Divya & Dinesh Peter, 2021). Quantum-inspired methods, which leverage phenomena such as quantum optimization algorithms, strive to maintain model performance by reducing computational requirements, thereby offering significant speedups for certain tasks (Pandey et al., 2023). Meanwhile, traditional model compression techniques reduce the size of neural networks by removing less important weights, sacrificing accuracy for lower latency (Francy & Singh, 2024). By integrating concepts from quantum mechanics into convolutional neural network (CNN) models, our approach seeks to address these limitations. We explore the potential of designing CNNs to balance improved inference times with minimal accuracy loss, creating a novel solution.
Within this context, we employ three key quantum-inspired principles: 1. quantum-inspired pruning: reducing model size by removing unnecessary parameters, guided by quantum approximation algorithms; 2. tensor decomposition: breaking down high-dimensional tensors into smaller components to reduce computational complexity; and 3. annealing-based matrix factorization: optimizing matrix factorization by using annealing techniques to find efficient representations of the data.
Our work addresses the following research question: How can principles from quantum computing be used to design and optimize CNNs to reduce latency and improve inference times, while still maintaining stable accuracies across various models?
In this paper, we propose a Quantum-Integrated Adaptive Networks (QIANets) – a comprehensive framework that integrates these quantum computing techniques into the DenseNet, GoogLeNet, and ResNet-18 architectures. To the best of our knowledge, this is the first attempt made to: 1) apply quantum computing-inspired algorithms into the models’ architectures to reduce computational requirements and achieve efficient performance improvements, and 2) specifically target these models.
The contributions of this work include:
  • QIANets: a comprehensive framework that integrates QAOA-inspired pruning, tensor decomposition and quantum annealing-inspired matrix factorization into three CNNs.
  • An exploration of the trade-offs between latency, inference time, and accuracy, highlighting the effects of applying quantum principles to CNN models for real-time optimization.
2 Related Works
Our proposed method builds upon the ideas of model compression & quantum-inspired techniques to improve the inference times of CNNs.
2.1 Model Compression Techniques:
Pruning is one of the most effective ways to accelerate CNNs. Cheng et al. (2018) provided a comprehensive review of model compression techniques for deep neural networks (DNNs), specifically focusing on parameter pruning methods: the removal of individual weights based on importance. This significantly reduces model size, while generally preserving the model’s performance.
Despite the advancements in parameter pruning, overall conventional pruning techniques have limitations: 1) when applied during the training phase, as such a method may be costly, and 2) the risk of prematurely removing important data. Hou et al. (2022) proposed a novel methodology called CHEX for training-based channel pruning and regrowing of channels throughout the training process. By employing a column subset selection (CSS) formulation, CHEX allocates and reassigns channels across layers, allowing for significant model compression without requiring a fully pre-trained model.
2.2 Quantum-Inspired Techniques for CNNs:
Quantum computing is currently recognized as a potential game-changer for various fields, including NLP, due to its ability to process complex data more efficiently than classical computers. Shi et al. (2021) proposed a quantum-inspired architecture for convolutional neural networks (QICNNs), using complex-valued weights to enhance the representational capacity of traditional CNNs. The authors display that their QICNNs achieve higher classification accuracy and faster convergence on datasets compared to standard CNNs. In contrast, our methodology prioritizes structural optimization for greater computational efficiency. We focus on reducing latency and improving inference times by employing various quantum techniques to decrease computational overhead in the tested CNNs.
Hu et al. (2022) set a high standard in the field by addressing the unique challenges of compressing quantum neural networks (QNNs). Their CompVQC framework leverages an alternating direction method of multipliers (ADMM) approach, achieving a remarkable reduction in circuit depth by over 2.5 times with less than 1% accuracy degradation. While their results in QNN compression are impressive, our research introduces a novel first-attempt technique that applies QAOA-inspired pruning, tensor decomposition and quantum annealing-inspired matrix factorization to classical CNNs. Our work can complement their approach, underscoring the potential of integrating quantum concepts to classical networks, and may lead to further improvements in model efficiency.
3 Methodology
3.1 Quantum-inspired pruning
Pruning is a widely applied technique for reducing CNN complexities; early studies have demonstrated the effectiveness of pruning in optimizing neural networks (LeCun et al., 1989; Hanson and Pratt, 1989; Hassibi et al., 1993). Our method implements a new approach to optimization through the Quantum Approximate Optimization Algorithm (QAOA). QAOA is designed to address problems in combinatorial optimization, which involves finding the best solution from a finite set of solutions.
Similarly, we adapt these principles by framing pruning as a probabilistic optimization problem. The goal is to identify the most important weights to retain while allowing others to approach zero. For a neural network layer represented by weights as a tensor: $W\in\mathbb{R}^{C_{out}\times C_{in}\times H\times W}$, we define the importance of each weight using its absolute value: $I_{i,j}=|W_{i,j}|$
To facilitate decision-making regarding weight retention, we normalize these importance scores with the softmax function to derive probabilities:
$P_{i,j}=\sum_{k,l}e^{I_{k,l}}e^{I_{i,j}}$
These probabilities are then utilized in a quantum-inspired decision-making process, where weights are selected for pruning based on a threshold $\mathit{\lambda}$, influenced by a hyperparameter known as layer sparsity.
$R_{i,j}=\begin{cases}1&\text{if }P_{i,j}\geq\lambda\\ 0&\text{otherwise}\end{cases}$
[…]
When implemented, we adopt an iterative approach for pruning the model across multiple stages. Each iteration recalculates the retain mask based on updated probabilities derived from the current weights. To enhance this process, we introduce a neighboring entanglement mechanism: when a weight is pruned, its adjacent weights in the tensor may also be pruned with a specified probability Pentangle. This mechanism simulates quantum entanglement, reflecting the idea that nearby weights may exhibit correlated behavior and can be pruned collectively. For convolutional layers specifically, this sequential pruning strategy is executed over several iterations, progressively reducing the number of parameters in the model while maintaining performance integrity.
3.2 Tensor Decomposition
Tensor decomposition further reduces the dimensionality of the weight tensor while preserving essential information for accurate predictions. This technique is inspired by Quantum Circuit Learning (QCL), where high-dimensional tensors are decomposed into lower-dimensional forms for efficient training of quantum circuits.
For a weight tensor $W\in\mathbb{R}^{C_{out}\times C_{in}\times H\times W}$, we use Singular Value Decomposition (SVD) on its flattened matrix representation $W_{f}\in\mathbb{R}^{C_{out}\times(C_{in}\cdot H\cdot W)}$, decomposing it as: $W_{f}=U\Sigma V^{T}$
Here, $U\in\mathbb{R}^{C_{out}\times r}$ and $V\in\mathbb{R}^{(C_{in}\cdot H\cdot W)\times r}$ are orthogonal matrices, and $\Sigma\in\mathbb{R}^{r\times r}$ is a diagonal matrix of singular values. The rank r, chosen as a hyperparameter, controls the compression level by retaining only the top r singular values, leading to a reduced rank approximation: $W_{f}\approx U_{r}\Sigma_{r}V_{r}^{T}$. This reduces the number of parameters and computational costs during inference.
After tensor decomposition, the original weight tensor is reconstructed using the truncated matrices, reshaping the compressed weights back into their original form. This process significantly decreases the number of parameters without greatly affecting model performance. We apply this method to each convolutional layer, creating lower-rank approximations to control the model’s capacity.
3.3 Quantum annealing-inspired matrix factorization
Quantum annealing solves optimization problems by evolving a system toward its lowest energy state. We apply this concept to factorize weight tensors, treating the factorization as an optimization problem aimed at minimizing the difference between the original weights and their factorized representation.
Given a weight matrix $W\in\mathbb{R}^{m\times n}$, we seek to factor it into two lower-dimensional matrices, $W_{1}\in\mathbb{R}^{m\times r}$and $W_{2}\in\mathbb{R}^{r\times n}$, where r is a hyperparameter that controls the rank: $W\approx W_{1}W_{2}$.
The objective is to minimize the reconstruction error: $L(W_{1},W_{2})=\|W-W_{1}W_{2}\|_{F}^{2}$
Here, $\|\cdot\|_{F}$ denotes the Frobenius norm, measuring the difference between the original and factorized matrices. We use an iterative optimization procedure inspired by quantum annealing to minimize loss.
The factorization employs gradient-based optimization, initializing $W_{1}$ and $W_{2}$ randomly. We iteratively minimize the loss function using an optimizer like LBFGS, suitable for small parameter sets and non-convex landscapes. This optimization simulates quantum annealing by gradually reducing the step size, ensuring convergence to a local minimum. Once complete, the compressed weight matrix is defined as: $W_{c}=W_{1}W_{2}$. This compressed matrix replaces the original matrix, reducing model complexity while maintaining performance.
Figure 1: An illustrative diagram showcasing the framework used for Quantum-Inspired Pruning, Tensor Decomposition, and Quantum Annealing-Inspired Matrix Factorization.
4 Experiments and results
We applied our method to compress three CNNs: DenseNet, GoogLeNet, and ResNet-18, all on the CIFAR-10 dataset. These networks were selected for their different design structures and computational demands, such as parameter count, depth, and layer types, providing a comprehensive assessment of our method’s effectiveness across different models. We evaluated the models for image classification performance, focusing on metrics such as inference time, speedup ratio, and accuracy.
Each experiment involved 1) applying the QIANets framework to the respective model architecture and 2) evaluating the models and comparing them to their baseline counterparts. The results, including networks’ changes before and after compression, are shown in Table 1.
Table 1: Model Performance Comparison Before and After Compression with rounded figures
Model BaseAccuracy BaseLatency New Accuracy NewLatency Compression Ratio
GoogleNet 94% 0.00096 T/I 86% 0.00083 T/I 1:0.52
ResNet 93% 0.000011 T/I 87% 0.00007 T/I 1:0.61
DenseNet 94% 0.000050 T/I 88% 0.00042 T/I 1:0.56
4.1 Experimental setup
The experiments were conducted within the cloud-based PyTorch framework, utilizing the CIFAR-10 dataset. The CIFAR-10 dataset, which consists of 60,000 32x32 RGB color images in 10 classes, with 6,000 images per class, was split into training and validation sets with an 80/20 ratio. To meet the input size requirements of the models, images were resized to 224x224 pixels. Data preprocessing included normalizing pixel values to the range [-1, 1] using the mean and standard deviation of the dataset. Moreover, data augmentation strategies (random horizontal flipping and random cropping) were applied to generate data variability and improve the models’ performance on unseen data.
All computations were accelerated using CUDA on an NVIDIA A40 GPU via Runpod. Each model underwent training for 50 epochs utilizing the Adam optimizer, with an initial learning rate of 0.001 and weight decay of 1e-4. To run our experiments, the compute used was approximately 1,615,680 TFLOPS-seconds. To ensure consistency across models, batch sizes of 128 were used for training, while batch sizes of 256 were employed for both evaluation and testing within the dataset.
4.2 Hyperparameter tuning
Hyperparameter tuning is performed using Optuna, a framework that implements a multitude of techniques to optimize certain parameters. Optuna experiments with various combinations of hyperparameters, including batch size, learning rate and ECA Kernel Size, and dynamically adjusts them based on each trial. Following each trial, validation accuracy is calculated to test the effectiveness of current parameters. This information dynamically refines the subsequent parameters, ultimately approaching optimized parameter configurations for model performance.
4.3 Model specific analysis
To effectively accommodate the unique architectures of each model, we made targeted adjustments to the QIANets method. These modifications were carefully designed to be minimal, ensuring that all models were trained and evaluated under consistent and fair conditions throughout the experiments.
4.3.1 GoogLeNet
GoogLeNet is a convolutional network with nine multi-scale processing Inception modules. Within our experiments, the QIANets framework targets the these modules, reducing the weight in their convolutions: layer sparsity of 0.1417 (only 14.17% of weights were pruned), while employing a rank of 41 to efficiently decompose and factorize these weights. ECA and Multi-Scale Fusion are applied to the outputs of the modules, integrating multi-scale attention into parallel branches.
Performance Progression Across Epochs: The training process involved 10 trials of 10 epochs each, followed by a final trial of 50 epochs. During Trial 0 of hyperparameter optimization, the model began with a validation accuracy of 21.08% at Epoch 1 and quickly progressed to 70.18% by Epoch 10, indicating rapid learning during the first stages of training. The highest validation accuracy result happened to be 80.19% throughout Trial 5. Ultimately, the final quantum-inspired GoogLeNet’s test accuracy was 86.65%, representing a notable improvement from the earlier 80.19%, and closely approaching the baseline accuracy of 94.29% after fine-tuning.
Loss Reduction: Over the 50 epochs, the model’s loss steadily decreased, showing consistent improvement. It began at 2.5205 in Epoch 1 trial 0 and reduced to 0.8256 by Epoch 50, effectively minimizing error throughout training. A key outcome of this experiment was the final 13.65% reduction in inference time, minimizing to 0.000835 seconds per image, which underscores the efficiency of our approach compared to the baseline GoogLeNet’s 0.000967 seconds.
4.3.2 DenseNet
We experimented on DenseNet – a CNN structured with 12 dense blocks, with layer-by-layer connections. This intricate connectivity requires careful application of QAOA pruning to ensure that weight removal does not disrupt the model’s residual stream and overall flow of information within the network. To enhance channel-wise interactions, we apply ECA and Multi-Scale Fusion after the dense blocks, allowing the model to leverage both local and global feature representations effectively.
Performance Progression Across Epochs: DenseNet was trained over 10 trials of 10 epochs each, followed by a final trial of 50 epochs to fine-tune performance. During Trial 0, the model began with a validation accuracy of 9.66% at Epoch 1 and exhibited minimal improvement by Epoch 10, reaching 10.34%. In contrast, Trial 1 demonstrated significant learning progression, starting at 27.53% and achieving a remarkable 81.33% by Epoch 10. The highest validation accuracy across all trials peaked at 86.65% in Trial 1, where the model achieved a layer sparsity of approximately 0.3779 (nearly 62% of the weights were pruned while maintaining performance). After extensive fine-tuning, the final quantum-inspired DenseNet achieved a test accuracy of 88.52%, a significant improvement from the earlier 86.65%, and approaching the baseline accuracy of 94.05%.
Loss Reduction: The model demonstrated steady loss reduction throughout the training process, beginning at 2.3028 during Epoch 1 in Trial 0 and decreasing to 0.5606 by Epoch 10 in Trial 1, indicating effective error minimization. This consistent decline reflects the model’s ability to optimize its parameters and improve performance across trials. One of the standout results of this experiment was the final reduction in inference time by 15.20%, dropping to 0.001043 seconds/image, marking a considerable improvement compared to the baseline DenseNet’s 0.00123 seconds.
4.3.3 ResNet-18
Lastly, ResNet-18 is a CNN characterized by its unique residual learning framework and shortcut connections that facilitate training in deep networks. Within our experiments, the QIANets framework targets the residual blocks in the model, reducing less significant weights and channels, detected by ECA’s straightforward 1D convolution. Feature maps are then combined by Multi-Scale Fusion and refined highlighting essential features across different scales.
Performance Progression: Throughout the trials, there were notable fluctuations in performance. The highest validation accuracy across all trials peaked at 91.42% in Trial 4, where the model achieved a layer sparsity of approximately 0.3779 (meaning nearly 62% of the weights were pruned while maintaining performance). After extensive fine-tuning, the final quantum-inspired ResNet-18 reached a test accuracy of 87.11%, a significant improvement from the earlier 84.56% (the highest accuracy before the final fine-tuning) and approaching the baseline accuracy of 93.11%.
Loss Reduction: The loss reduction across trials also followed a clear downward trend. In Trial 3, with a layer sparsity of 0.4805 and a rank of 10, the validation loss dropped from 1.9847 in the first epoch to 0.6321 by the tenth epoch, indicating better model convergence. One of the standout results of this experiment was the final reduction in inference time by 36.4%, dropping to 0.00007 seconds/image, marking a improvement compared to the baseline ResNet’s 0.00011 seconds/image.
5 Conclusion
In this paper, we introduced the QIANets framework and applied it to three prominent convolutional neural networks—DenseNet, GoogLeNet, and ResNet—with the objective of reducing latency and improving inference times while maintaining minimal accuracy loss. Our experimental results showed varying degrees of success when applying quantum-inspired techniques, revealing important insights into the trade-offs between latency and accuracy across different architectures.
6 Limitations
While our results demonstrate the potential of QIANets and quantum-inspired principles in model compression, they also reflect on several factors that influence the performance of our approach:
  1. 1.
    Data Constraints: The evaluation was restricted to the relatively simple CIFAR-10 dataset, which may represent the full diversity, complexity, or scalability challenges encountered in real-world scenarios that larger datasets may exhibit. Additionally, due to the computational expense of each run, the approach was only tested on a limited number of trials.
  2. 2.
    Model Adaptation: The lack of adaptation across different model architectures may hinder the QIANets framework’s ability to optimize the balance between latency and accuracy. Performance in certain cases does not guarantee similar results across architectures without mandatory model-specific adjustments, which can complicate future adaptations.
  3. 3.
    Hardware Limitations: This study does not consider hardware-specific limitations. Our techniques have not yet been optimized for specialized hardware, such as custom FPGAs or GPUs, which could potentially reduce latency and improve data throughput.
  4. 4.
    Scope of Focus: The narrow focus on CNNs excludes newer, more advanced architectures, restricting the scope of our framework’s potential. Integrating our concept with transformers and other attention-based models could significantly expand QIANets’ applicability.
Future works should address these limitations by conducting more in-depth experiments that assess both the scalability and practical relevance of the quantum-inspired techniques.
Acknowledgments and Disclosure of Funding
Use unnumbered first level headings for the acknowledgments. All acknowledgments go at the end of the paper before the list of references. Moreover, you are required to declare funding (financial activities supporting the submitted work) and competing interests (related financial activities outside the submitted work). More information about this disclosure can be found at: https://neurips.cc/Conferences/2024/PaperInformation/FundingDisclosure.
Do not include this section in the anonymized submission, only in the final paper. You can use the ack environment provided in the style file to automatically hide this section in the anonymized submission.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The introduction clearly introduces QIANets and how the method aims to integrate the quantum-inspired strategies, accurately reflecting the paper’s scope. The claims made in the abstract/introduction are supported by the results obtained.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: The paper discusses limitations in section 5, explaining the data and hardware constraints and limited number of runs. We state the lack of adaptation and provide avenues for future work to experiment on a more expansive range of models and datasets.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory Assumptions and Proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: The paper does not include any theoretical results.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental Result Reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: The code and instructions on how to reproduce the experiments are included in the supplementary material. Within the paper, the model is described in detail with code snippets and is reproducible using the information provided.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: Code is provided with instructions as supplemental material.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental Setting/Details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: The paper covers the experimental set up, including training and test details in section 4.1. Full details are included in the code.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment Statistical Significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [No]
  34. Justification: Statistical significance tests are not included as we could not conduct multiple trails of our finalized method due to the computational expense. As a result, error bars and confidence intervals are not included in the results.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments Compute Resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: The paper specifies our use of an NVIDIA A40 GPU and the approximate amount of compute used in section 4.1.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code Of Ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: The research conducted complies fully with all of the ethical guidelines listed.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader Impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [N/A]
  49. Justification: The research has no potential societal impacts besides increasing the speed of responses from preexisting models.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: The paper does not introduce/include data or models with such risks.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: The paper properly cites the CIFAR-10 dataset that the models were tested on.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New Assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: All the code is provided as supplemental material, and there are no new assets besides the proposed method.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and Research with Human Subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.